OpenStudy (anonymous):

How do I solve x^3-4x?

jimthompson5910 (jim_thompson5910):

you want to factor?

OpenStudy (anonymous):

Yes, its equal to 0

jimthompson5910 (jim_thompson5910):

x^3-4x = 0 x(x^2 - 4) = 0 ... factor out x x(x-2)(x+2) = 0 ... use the difference of squares rule

jimthompson5910 (jim_thompson5910):

You can then break up x(x-2)(x+2) = 0 into x = 0 or x-2 = 0 or x+2 = 0 basically you set each factor equal to zero

jimthompson5910 (jim_thompson5910):

now solve these 3 equations: x = 0 or x-2 = 0 or x+2 = 0 well one is already solved for you

OpenStudy (anonymous):

Thank you so much, I didn't see that.

OpenStudy (anonymous):

can you help me with one more please?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

okay, x^4-1

jimthompson5910 (jim_thompson5910):

x^4 - 1 = x^4 - 1^2 use the difference of squares rule to get ???

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

That one is equal to zero also.

jimthompson5910 (jim_thompson5910):

ok try thinking of it like this (x^2)^2 - 1^2 if you factored that, what would you get?

OpenStudy (anonymous):

why is it 1^2?

jimthompson5910 (jim_thompson5910):

because 1 = 1^2

jimthompson5910 (jim_thompson5910):

I'm formatting the given expression as a difference of squares

jimthompson5910 (jim_thompson5910):

so you can factor using that rule

OpenStudy (anonymous):

so how do I solve it? I'm sorry, I'm just really confused.

jimthompson5910 (jim_thompson5910):

say you had x^2 - 1 how would you factor that?

OpenStudy (anonymous):

(x-1)(x+1) ? right?

jimthompson5910 (jim_thompson5910):

so (x^2)^2 - 1 factors to what?

OpenStudy (anonymous):

I don't know.

jimthompson5910 (jim_thompson5910):

let z = x^2

jimthompson5910 (jim_thompson5910):

z^2 - 1 factors to what?

OpenStudy (anonymous):

x^4 - 1

OpenStudy (anonymous):

so what does that haveto do with solving it?

jimthompson5910 (jim_thompson5910):

I'm helping you factor x^4 - 1

jimthompson5910 (jim_thompson5910):

z^2 - 1 = (z+1)(z-1) replace every z with (x^2) z^2 - 1 = (z+1)(z-1) (x^2)^2 - 1 = ((x^2)+1)((x^2)-1) (x^2)^2 - 1 = (x^2+1)(x^2-1) now factor x^2-1 to get (x-1)(x+1) (x^2)^2 - 1 = (x^2+1)(x-1)(x+1)